OpenStudy (anonymous):

Write a third-degree polynomial function whose zeros are 1, −3, and 4.

jimthompson5910 (jim_thompson5910):

The zeros are 1, −3, and 4, so x = 1, x = -3, x = 4 x - 1 = 0, x + 3 = 0, x - 4 = 0 (x - 1)(x + 3)(x - 4) = 0 Do you know what the next step is?

OpenStudy (anonymous):

im so confused

jimthompson5910 (jim_thompson5910):

do you see how I got that so far?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

the next step is to expand the left side

jimthompson5910 (jim_thompson5910):

we use the distributive property to do that

jimthompson5910 (jim_thompson5910):

(x - 1)(x + 3)(x - 4) = 0 (x - 1)[ x(x - 4) + 3(x - 4) ] = 0 (x - 1)[ x^2 - 4x + 3x - 12 ] = 0 (x - 1)( x^2 - x - 12 ) = 0 x( x^2 - x - 12 ) - 1( x^2 - x - 12 ) = 0 x^3 - x^2 - 12x - x^2 + x + 12 = 0 x^3 - 2x^2 - 11x + 12 = 0

jimthompson5910 (jim_thompson5910):

So all of that shows you how (x - 1)(x + 3)(x - 4) turns into x^3 - 2x^2 - 11x + 12

OpenStudy (anonymous):

im sort of getting it

jimthompson5910 (jim_thompson5910):

keep practicing it and I'm sure it'll click more and more

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thank you

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

Is there a polynomial function with any given number of zeros? What is its degree?

jimthompson5910 (jim_thompson5910):

what do you mean?

OpenStudy (anonymous):

im staring at the question right now trying to understand it myself

jimthompson5910 (jim_thompson5910):

oh, if you have 3 zeros, then the degree will be 3

jimthompson5910 (jim_thompson5910):

if there were 4 zeros, then the degree is 4

OpenStudy (anonymous):

ok thanks

jimthompson5910 (jim_thompson5910):

yw